# Conformal online model aggregation

### v1

@@ 1 Introduction
Conformal prediction has emerged as a significant approach within the field of machine learning, offering a reliable procedure for uncertainty quantification. Unlike traditional methods that rely on strict distributional assumptions, conformal prediction provides a flexible framework by transforming point predictions obtained with any black-box model into confidence sets with valid finite-sample marginal coverage. Initially introduced by Vovk et al., (2005), conformal prediction has gained increasing prominence and has been extended beyond the i.i.d. (or exchangeable) data setting, some examples are Tibshirani et al., (2019); Barber et al., (2023); Gibbs and Candès, (2021).
A significant challenge in the application of conformal prediction arises in the context of model selection or aggregation. In fact, the conventional approach involves fixing the underlying machine learning (or prediction) model in advance and then obtaining an interval with a pre-specified level of marginal coverage. Furthermore, since with different models it is possible to obtain sets with the same theoretical guarantees, it becomes difficult to assess which model is the most suitable.
@@ 1.1 Problem setup
We focus our attention on an online setting in which data arrive sequentially over time. A plausible scenario arises when dealing with a continuous flow of data; in symbols, for each iteration $t=1,2,...,T$, the covariate-response pair $z^{(t)}=(x^{(t)},y^{(t)})\in\mathcal{X}\times\mathcal{Y}\subseteq\mathbb{R}^{d}\times\mathbb{R}$ is observed.
In our scenario, we assume the presence of $K$ different prediction algorithms (or experts), denoted as $\hat{\mu}_{k}$, used to predict $y^{(t)}$ based on a given set of covariates $x^{(t)}$. As a result, in each iteration, we can derive $K$ distinct conformal prediction intervals (one for each model) $\mathcal{C}_{1}^{(t)},\dots,\mathcal{C}_{K}^{(t)}$ using the information from the set $\{z^{(i)}\}_{i=1}^{t-1}$. Actually, one can choose to use only a subset of $\{z^{(i)}\}_{i=1}^{t-1}$ to produce the interval due to computational issues or distribution shifts. If the data exhibit exchangeability, these sets have valid marginal coverage. This implies that, for any $\alpha\in(0,1)$, we achieve:
@@
In light of the fact that our method is designed to return a single set from a collection of intervals, the initial question that arises is how to merge efficiently different conformal prediction intervals. It is crucial to highlight that these intervals are not independent, as they are constructed using the same data. One potential solution is the approach proposed by Gasparin and Ramdas, (2024), which involves a weighted majority vote procedure. Let $w^{(t)}=(w^{(t)}_{1},\dots,w^{(t)}_{K})$ be a set of weights, such that
$w^{(t)}_{k}\in[0,1],\quad k=1,\dots,K,$ (2)
$\sum_{k=1}^{K}w^{(t)}_{k}=1,$ (3)
@@
$\mathcal{C}^{(t)}:=\left\{y\in\mathcal{Y}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y\in\mathcal{C}_{k}^{(t)}\right\}>\frac{1+u}{2}\right\},$ (4)
@@
where $u$ is a realization from a uniformly distributed random variable in the interval $[0,1]$. If the weights are chosen a priori, it is possible to prove that if the starting intervals respect the property (1) then $\mathcal{C}^{(t)}$ has a valid $1-2\alpha$ coverage. In light of the selection of weights based on the data, they themselves become random quantities. Specifically, we have that $w^{(t)}$ is a realization of a random vector $W^{(t)}$ mapping the first $t-1$ observations into the $(K-1)$-dimensional simplex. Formally, we have $W^{(t)}:(\mathcal{X}\times\mathcal{Y})^{t-1}\to[0,1]^{K}$ with the additional constraint that the weights have to sum to one. In the following, we define the miscoverage indicator for the $k$-th interval at time $t$ as $\phi_{k}^{(t)}:=\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\}$.
Proposition 1.1.
Let $\mathcal{C}_{1}^{(t)},\dots,\mathcal{C}_{K}^{(t)}$ be $K\geq 2$ different conformal prediction intervals that satisfy the property in (1). Let $W^{(t)}$ be a random vector having values in the $(K-1)$-dimensional simplex. Under the assumption that $\phi_{k}^{(t)}$ and $W_{k}^{(t)}$ are independent or negatively associated, for all $k=1,\dots,K$, then the set defined in (4) is such that $\mathbb{P}(Y^{(t)}\in\mathcal{C}^{(t)})\geq 1-2\alpha$.
@@
Due to the property described in (1) we have that $\mathbb{E}[\phi_{k}^{(t)}]\leq\alpha$, $k=1,\dots,K$. Now we can write
@@
$\begin{split}\mathbb{P}\left(Y^{(t)}\notin\mathcal{C}^{(t)}\right)&=\mathbb{P}\left(\sum_{k=1}^{K}W_{k}^{(t)}\phi_{k}^{(t)}\geq\frac{1}{2}-U/2\right)\stackrel{{\scriptstyle(i)}}{{\leq}}2\mathbb{E}\left[\sum_{k=1}^{K}W_{k}^{(t)}\phi_{k}^{(t)}\right]\\ &\stackrel{{\scriptstyle(ii)}}{{\leq}}2\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]\,\mathbb{E}[\phi_{k}^{(t)}]\leq 2\alpha\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]=2\alpha,\end{split}$
where $(i)$ is valid due to the randomized Markov’s inequality described in Ramdas and Manole, (2023), while $(ii)$ holds due to the assumption regarding the independence or the negative association between the variables $\phi_{k}^{(t)}$ and $W_{k}^{(t)}$. In particular, $(ii)$ holds with equality under independence. ∎
The case of independence arises when the data are i.i.d. and the $K$ models are pre-trained and not iteratively updated or, more generally, when the models are trained in a different dataset than the one used for weights.
We remark the fact that if randomization is not allowed, then one can set the value of $u$ to 00 and the coverage guarantee still holds since $u>0$ almost surely. In addition, we highlight that if the weight vector takes the form $w^{(t)}_{k}=1$, and $w^{(t)}_{j}=0$ for all $j\neq k$, then our procedure reduces to the selection of the $k$-th model. A potential drawback of the procedure is that, even when initiated with a collection of intervals, it may produce a union of intervals. This does not occur if the intersection of the initial intervals is non-empty.
Consistent with our previous discussion, the loss of the different models is measured through the size of the outputted intervals. We define the loss function $\ell(\cdot)$ as
$\ell:=\ell(\mathcal{C})=\text{Lebesgue measure}(\mathcal{C}),$ (5)
where $\mathcal{C}$ represent an arbitrary set. In general, one could choose some nondecreasing function of the Lebesgue measure, but the identity seems to be a sensible canonical choice, so we stick to it in this paper. At each iteration, we observe losses of the experts $\ell^{(t)}=(\ell_{1}^{(t)},\dots,\ell_{K}^{(t)})\in\{\mathbb{R}^{+}\}^{K}$ and the cumulative loss for a given method after $t$ rounds is given by $L_{k}^{(t)}=\ell^{(1)}_{k}+\dots+\ell^{(t)}_{k}$.
A notable property of the majority vote is that it generates sets that are never larger than twice the weighted average size of the initial sets. For completeness, we present here an useful Lemma, which will be used in the following.
Lemma 1.2.
Let $\mathcal{C}^{(t)}$ be the set defined in (4), $\ell_{M}^{(t)}=\ell(\mathcal{C}^{(t)})$ be the Lebesgue measure associated to $\mathcal{C}^{(t)}$ and $w^{(t)}$ be a realization of $W^{(t)}$. Then,
The proof is given in Gasparin and Ramdas, (2024).
@@
From equation (6), it is now clear that our goal is to find a good weighting system in order to narrow down the size of $\mathcal{C}^{(t)}$, particularly we aim to assign a higher weight to experts that consistently yield smaller intervals compared to others.
@@
In the following, we describe a general methodology to infer the weights assigned to the various experts over time. In particular, we start by describing the method under an i.i.d. setting and then extend the approach under a (possible) continuous distribution shift over time, as described in Gibbs and Candès, (2021). We will then demonstrate the practical effectiveness of the method through a series of simulations and examples using real-world data.
2 Related work
The so-called full conformal prediction was initially introduced by Vovk et al., (2005), while more computationally efficient versions were proposed by Papadopoulos et al., (2002), Lei et al., (2018) and Romano et al., (2019). In particular, the latter methods propose dividing the training set into two parts: a proper training set and a calibration set. All these approaches provide reliable uncertainty quantification for a single black-box algorithm and are grounded on the assumption that observations are independent and identically distributed or at least exchangeable.
Lastly, outside the conformal framework, online learning is used to develop sequential prediction algorithms when multiple agents construct a prediction at each time. In this particular setting, it is customary that at each iteration a weight is assigned to the agent, which is subsequently updated in accordance with its performance. One of the most widely employed methods is the Exponential Weighted (EW) algorithm (Freund and Schapire,, 1997), which updates the weights in an inversely proportional manner to the exponential of the loss incurred at the $t$-th iteration. One of its extensions is proposed in de Rooij et al., (2014), and it has been demonstrated to be effective even in scenarios where the data is not adversarial.
@@ 3 Dynamic merging via exponential weighted majority
As illustrated in Section 1.1, our goal is to achieve a set that is as small as possible, particularly (6) asserts that this can be accomplished by learning a weight system that prioritizes methods yielding smaller intervals. Furthermore, we want that the cumulative loss of our proposed method, indicated by $L_{M}^{(t)}=\ell_{M}^{(1)}+\dots+\ell_{M}^{(t)}$, is not too large if compared to the cumulative loss incurred by the best expert up to time, indicated by
$L_{*}^{(t)}:=\min_{k\in\{1,\dots,K\}}L_{k}^{(t)}.$
@@
$w_{k}^{(1)}\leftarrow 1/K,L_{k}^{(0)}\leftarrow 0,$ $k=1,\dots,K$;
for rounds $t=1,\dots,T$ do do
       $\mathcal{C}^{(t)}\leftarrow\left\{s\in\mathcal{S}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\{s\in\mathcal{C}_{k}^{(t)}\}>\frac{1}{2}\right\}$;
@@
Our proposed solution is to employ the AdaHedge algorithm, which dynamically adjusts the learning parameter over time as follows. The weights at each round are assigned according to $w^{(t)}_{k}\propto\exp\{-\eta^{(t)}L^{(t-1)}_{k}\}$, $k=1,\dots,K$, with
@@
It is possible to bound the total measure of the sets produced by the majority vote method.
Proposition 3.1.
The total Lebesgue measure of the weighted majority vote set over $T$ rounds can be bounded as
@@
As explained at the beginning, a relatively open question in the conformal prediction literature is how to do model selection or aggregation using conformal prediction: if we have $K$ different prediction algorithms that one could be using within conformal prediction, how do we combine their predictions so that we can essentially do as well as the “best” predictor? Since all methods do provide the same coverage level of $1-\alpha$ under i.i.d. data, we use a working definition of the “best” predictor as the one whose expected interval length is shortest.
Below, we provide one answer to this question in an online or streaming data setting, by employing the algorithms introduced in the previous section. When our dynamic merging Algorithm 1 is used (with constant, adaptive or another stepsize rule) in this conformal context, we title the method as “conformal online model aggregation”, or COMA for short. In the i.i.d. setting, it is often the case that our adaptive stepsize update rule results in the weight vector rapidly putting almost full (unit) weight on the best predictor, resulting effectively in online model “selection”, but experiments in non-i.i.d. settings suggest that the weight vector can meaningfully fluctuate between different predictors if they perform better in different periods.
We start by examining our proposed methods, in a scenario where data are i.i.d., this implies that each algorithm at each iteration provides an interval with the guarantee described in (1), while our procedure has a valid $1-2\alpha$ coverage level.
To examine the practical behavior of COMA, we apply it to a real-world dataset. Specifically, the dataset comprises $75000$ observations pertaining to Airbnb apartments in New York City. The response variable is represented by the logarithm of the nightly price, while the covariates include information about the apartment and its geographical location. The data has been partitioned into 75 rounds, each consisting of $1000$ observations. At each iteration, $K$ split-conformal prediction intervals are constructed using various algorithms. In the first simulation, $K=4$ regression algorithms were employed: linear model, lasso, ridge, and random forest. In the second scenario $K=5$ and also a neural network was utilized.
@@
The weights have been updated with the COMA with fixed or adaptive $\eta$ methods described above. For Algorithm 1 with variable learning parameter throughout the iterations, two different values were selected, $\eta=0.1$ and $\eta=1$. In Figure 1, the hedge loss during various iterations in both scenarios is presented, along with the loss obtained by the best performing method(s). It can be observed that in both cases, the hedge loss of the COMA algorithm with $\eta=1$ and of the adaptive COMA stabilizes at the level of loss obtained by the random forest after a certain number of iterations, which is the best method. On the other hand, if $\eta$ is set to 0.1 the loss exhibits a slower decrease, attributed to the lower value of the learning parameter. In Figure 6 and Figure 5 of Appendix A, we report the weights assumed by the different regression algorithms during certain iterations. It can be seen that the COMA with adaptive parameter, after a few iterations, concentrates all the mass in the random forest, while the COMA algorithm with fixed learning parameter is more conservative in assigning the weights. This fact can be explained by the plot in Figure 7 of Appendix A where it is possible to note that in both cases ($K=4$ and $K=5$) the learning parameter $\eta^{(t)}$ is always higher than the levels $\eta=1$ and $\eta=0.1$.

### v2

@@ 1 Introduction
Conformal prediction has emerged as a significant approach within the field of machine learning, offering a reliable procedure for uncertainty quantification. Unlike traditional methods that rely on strict distributional assumptions, conformal prediction provides a flexible framework by transforming point predictions obtained with any black-box model into prediction sets with valid finite-sample marginal coverage. Initially introduced by Vovk et al., (2005), conformal prediction has gained increasing prominence and has been extended beyond the i.i.d. (or exchangeable) data setting, some examples are Tibshirani et al., (2019); Barber et al., (2023); Gibbs and Candès, (2021).
A significant challenge in the application of conformal prediction arises in the context of model selection or aggregation. The conventional approach involves fixing the underlying machine learning (or prediction) model in advance and then obtaining an interval with a pre-specified level of marginal coverage. Since different models yield sets with the same theoretical coverage guarantees, the natural goal is to select the best, or perform as well as the best, of a collection of models. In this work, “best” refers to the set with smallest size, such as cardinality for a discrete set, or Lebesgue measure for a continuous set.
@@ 1.1 Problem setup
We focus our attention on an online setting in which data arrive sequentially over time. A plausible scenario arises when dealing with a continuous flow of data; in symbols, for each iteration $t=1,2,...,T$, a covariate-response pair $z^{(t)}=(x^{(t)},y^{(t)})\in\mathcal{X}\times\mathcal{Y}$ representing an instance from the random vector $Z^{(t)}=(X^{(t)},Y^{(t)})$ is observed.
In our scenario, we assume the presence of $K$ different prediction algorithms (or experts), denoted as $\hat{\mu}_{k}$, used to predict $Y^{(t)}$ based on a given set of covariates $X^{(t)}$. As a result, in each iteration, we can derive $K$ distinct conformal prediction sets (one for each model) $\mathcal{C}_{1}^{(t)},\dots,\mathcal{C}_{K}^{(t)}$ based on the information from ($(\{z^{(i)}\}_{i=1}^{t-1},x^{(t)})$. Actually, one can choose to use only a subset of $\{z^{(i)}\}_{i=1}^{t-1}$ to produce the interval due to computational issues or distribution shifts. If the data exhibit exchangeability, these sets have valid marginal coverage. This implies that, for any $\alpha\in(0,1)$, we achieve:
@@
In light of the fact that our method is designed to return a single set from a collection of intervals, the initial question that arises is how to efficiently merge different conformal prediction sets (i.e. conformal prediction sets produced by different underlying prediction algorithms). These intervals are not independent, as they are constructed using the same data. One potential solution is an approach recently proposed by Gasparin and Ramdas, (2024), which involves a weighted majority vote procedure. Let $w^{(t)}=(w^{(t)}_{1},\dots,w^{(t)}_{K})$ be a set of weights, such that
$w^{(t)}_{k}\in[0,1],\quad k=1,\dots,K,\quad\text{ and }\sum_{k=1}^{K}w^{(t)}_{k}=1,$ (2)
@@
$\mathcal{C}^{(t)}:=\left\{y\in\mathcal{Y}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y\in\mathcal{C}_{k}^{(t)}\right\}>\frac{1+u^{(t)}}{2}\right\},$ (3)
@@
where $u^{(t)}$ is a realization from a uniformly distributed random variable in the interval $[0,1]$ independent of all the data. If the weights are chosen a priori (independent of the data), Gasparin and Ramdas, (2024) prove that if the starting intervals are marginally valid (they respect property (1)) then $\mathcal{C}^{(t)}$ has $1-2\alpha$ coverage, and the factor 2 is unavoidable in general.
There are two hurdles to applying their ideas directly. First, in light of the selection of weights based on the data, the weights themselves become random quantities, so their theory does not immediately apply (especially for non-i.i.d. data). The second question is how one should update the weights in a theoretically sound way that performs well practically. The rest of this paper will present ways to overcome these hurdles.
1.2 Related work
The so-called full conformal prediction was initially introduced by Vovk et al., (2005), while more computationally efficient versions were proposed by Papadopoulos et al., (2002), Lei et al., (2018), Romano et al., (2019) and Romano et al., (2020). In particular, the latter methods propose dividing the training set into two parts: a proper training set and a calibration set. All these approaches provide reliable uncertainty quantification for a single black-box algorithm and are grounded on the assumption that observations are independent and identically distributed or at least exchangeable. Humbert et al., (2023) proposes a method for aggregating the results of various agents within a conformal prediction framework. This method specifically addresses the context of federated learning, where privacy concerns restrict agent operations. The problem of model selection, within the realm of conformal prediction, is studied by Yang and Kuchibhotla, (2021) and Fan et al., (2023). Our work is distinct from theirs, as it operates within an online setting and can be extended to scenarios where data are not independently and identically distributed. Furthermore, rather than performing model selection, our approach focuses on model aggregation.
Lastly, outside the conformal framework, online learning is used to develop sequential prediction algorithms when multiple agents construct a prediction at each time. In this particular setting, it is customary that at each iteration a weight is assigned to the agent, which is subsequently updated in accordance with its performance. One of the most widely employed methods is the Exponential Weighted (EW) algorithm (Freund and Schapire,, 1997), which updates the weights in an inversely proportional manner to the exponential of the loss incurred at the $t$-th iteration. See Cesa-Bianchi and Lugosi, (2006) for an introduction of the method. One of its extensions is proposed in de Rooij et al., (2014), and it has been demonstrated to be effective even in scenarios where the data is not adversarial.
1.3 Summary of contributions
We introduce a method for updating the weights in (3) based on the performance obtained by the different models; where the performance is measured through the dimension of the set (cardinality if $\mathcal{Y}$ is a discrete set or size if $\mathcal{Y}\subseteq\mathbb{R}$). This method relies on the exponential weighted algorithm and enables the establishment of an upper bound on the observed loss in relation to the best-performing algorithm, after $T$ iterations. In addition, we provide the theoretical assumptions under which our method guarantees a valid prediction set of level $1-2\alpha$. We will then demonstrate the practical effectiveness of the method through a series of simulations and examples using real-world data.
2 Weighted majority with data-dependent weights
Suppose we have that $w^{(t)}$ is a realization of a random vector $W^{(t)}$ that maps the first $t-1$ observations of $\{Z^{(i)}\}_{i=1}^{t-1}$ into the $(K-1)$-dimensional simplex $\Delta_{K-1}:=\{w\in[0,1]^{K}:\mathbf{1}^{T}w=1\}$, or more succinctly, $W^{(t)}:(\mathcal{X}\times\mathcal{Y})^{t-1}\to\Delta_{K-1}$. In the following, we define the miscoverage indicator for the $k$-th set at time $t$ as $\phi_{k}^{(t)}:=\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\}$ and $[K]:=\{1,\dots,K\}$. We will make one of the following assumptions:
  1. 1.
    Conditional coverage:
    $\mathbb{E}\left[\phi_{k}^{(t)}\mid\{Z^{(i)}\}_{i=1}^{t-1}\right]\leq\alpha,\text{ for all }k\in[K].$ (4)
    This holds, for example, for conformal prediction when the data are i.i.d.
  2. 2.
    Marginal coverage and negative total elementwise correlation between $\phi^{(t)}$, $W^{(t)}$:
    $\mathbb{E}\left[\phi_{k}^{(t)}\right]\leq\alpha\text{ for }k\in[K],\leavevmode\nobreak\ \text{ and }\mathbb{E}\left[\sum_{k=1}^{K}\left(\phi_{k}^{(t)}-\mathbb{E}[\phi_{k}^{(t)}]\right)\left(W_{k}^{(t)}-\mathbb{E}[W_{k}^{(t)}]\right)\right]\leq 0.$ (5)
    $\phi^{(t)}$, $W^{(t)}$ could be negatively correlated if, for example, it happens to be the case that predictors that were assigned high weights (based on past data) are less likely to make mistakes.
Theorem 2.1.
Let $\mathcal{C}_{1}^{(t)},\dots,\mathcal{C}_{K}^{(t)}$ be $K\geq 2$ different prediction intervals for $Y^{(t)}$, let $W^{(t)}$ be a random vector in $\Delta_{K-1}$ depending only on $\{Z^{(i)}\}_{i=1}^{t-1}$, and consider the weighted majority vote set $\mathcal{C}^{(t)}$ in (3). It always holds that $\mathbb{P}(Y^{(t)}\notin\mathcal{C}^{(t)})\leq 2\mathbb{E}[\max_{k}(\phi_{1}^{(t)},\dots,\phi^{(t)}_{K})]$. If either assumption (4) or (5) holds, then $\mathbb{P}(Y^{(t)}\notin\mathcal{C}^{(t)})\leq 2\alpha.$
The proof of the theorem is given in Appendix C.
Remark 2.2.
The first part of the theorem states, without any assumption, that the probability of miscoverage is bounded by $2\mathbb{E}[\max_{k}(\phi_{1}^{(t)},\dots,\phi^{(t)}_{K})]$. This can translate into a trivial bound (like 1) if there is at least one algorithm that always makes an error, while it can be approximately $2\alpha$ in the case where miscoverage events are strongly correlated and $\mathbb{E}[\phi^{(t)}_{k}]\leq\alpha$, for each $k=1,\dots,K$.
Remark 2.3.
An elementary instance where the negative total elementwise correlation stated in Assumption 2 in Theorem 2.1 is valid can be described as follows. Consider the scenario involving just two models, resulting in a weight vector denoted by $(W_{1},1-W_{1})$, where the superscript $(t)$ is excluded for simplicity. Moreover, assume $\phi_{1}=\phi_{2}=\phi$, which might occur when two algorithms deliver nearly identical results, with one slightly outperforming the other. If $\mathrm{cov}(W_{1},\phi)=\rho$, by definition it follows that $\mathrm{cov}(1-W_{1},\phi)=-\rho$; this implies that $\sum_{k=1}^{2}\mathbb{E}[\phi W_{k}]=\sum_{k=1}^{2}\mathbb{E}[\phi]\,\mathbb{E}[W_{k}]$. Note that this is merely a basic illustration, and the assumption is explored in greater depth in Appendix B via simulations and real data examples.
We highlight that if one prefers to avoid randomization, then one can set the value of $u^{(t)}$ to 00 and the coverage guarantee still holds. In addition, we note that if the weight vector takes the form $w^{(t)}_{k}=1$, and $w^{(t)}_{j}=0$ for all $j\neq k$, then our procedure reduces to the selection of the $k$-th model. A potential drawback of the procedure is that it may produce a union of intervals, even when the input sets are intervals. This does not occur if the intersection of the initial intervals is non-empty (Gasparin and Ramdas,, 2024).
Building on our earlier discussion, consider two canonical examples for the loss as follows: for regression tasks, the loss is quantified by the interval’s size, and for classification tasks, by the count of elements within the set. We define the loss function $\ell(\cdot)$ as
$\ell:=\ell(\mathcal{C})=\begin{cases}\text{Lebesgue measure}(\mathcal{C}),\quad&\mathrm{if}\leavevmode\nobreak\ \mathcal{Y}\subseteq\mathbb{R},\\ |\mathcal{C}|,\quad&\mathrm{if}\leavevmode\nobreak\ \mathcal{Y}=\{y_{1},\dots,y_{D}\},\end{cases}$ (6)
where $\mathcal{C}$ denotes any set, $|\mathcal{C}|$ represents the cardinality of $\mathcal{C}$, and $\{y_{1},\dots,y_{D}\}$ describes a discrete set with $D$ elements. Specifically, (6) varies when dealing with a regression or classification problem. In general, one could choose some nondecreasing function of the Lebesgue measure (or of the cardinality), but the identity seems to be a sensible canonical choice, so we stick to it in this paper. At each iteration, we observe losses of the experts $\ell^{(t)}=(\ell_{1}^{(t)},\dots,\ell_{K}^{(t)})\in\{\mathbb{R}^{+}\}^{K}$ and the cumulative loss for a given method after $t$ rounds is given by $L_{k}^{(t)}:=\ell^{(1)}_{k}+\dots+\ell^{(t)}_{k}$.
A notable property of the majority vote is that it generates sets that are never larger than twice the weighted average size of the initial sets. We present here a useful result, which will be used in the following.
Lemma 2.4.
Let $\mathcal{C}^{(t)}$ be the set defined in (3), $\ell_{M}^{(t)}=\ell(\mathcal{C}^{(t)})$ be the loss function defined in (6) associated to $\mathcal{C}^{(t)}$ and $w^{(t)}$ be any realization of $W^{(t)}$. Then,
@@ 2 Weighted majority with data-dependent weights
We start with the discrete case $\mathcal{Y}=\{y_{1},\dots,y_{D}\}$. In particular, we have that
@@
$\begin{split}\ell_{M}^{(t)}&=\sum_{d=1}^{D}\mathds{1}\left\{\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y_{d}\in\mathcal{C}_{k}^{(t)}\right\}>\frac{1}{2}+\frac{u^{(t)}}{2}\right\}\leq\sum_{d=1}^{D}\mathds{1}\left\{\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y_{d}\in\mathcal{C}_{k}^{(t)}\right\}>\frac{1}{2}\right\}\\ &\stackrel{{\scriptstyle(i)}}{{\leq}}\sum_{d=1}^{D}2\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y_{d}\in\mathcal{C}_{k}^{(t)}\right\}=2\sum_{k=1}^{K}w_{k}^{(t)}\sum_{d=1}^{D}\mathds{1}\left\{y_{d}\in\mathcal{C}_{k}^{(t)}\right\}=2\sum_{k=1}^{K}w_{k}^{(t)}\ell(\mathcal{C}_{k}^{(t)}),\end{split}$
where $(i)$ holds due to the fact that $\mathds{1}\{x>1\}\leq x$, if $x\geq 0$. The proof in the case $\mathcal{Y}\subseteq\mathbb{R}$ follows the same strategy, it simply replaces the summation in $d$ by an integral and is given in Gasparin and Ramdas, (2024). ∎
@@
From (7), it is now clear that our goal is to find a good weighting system in order to narrow down the size of $\mathcal{C}^{(t)}$, particularly we aim to assign a higher weight to experts that consistently yield smaller intervals compared to others.
@@
In the following, we describe a general methodology to infer the weights assigned to the various experts over time. In particular, we start by describing the method under an i.i.d. setting and then extend the approach under a (possible) continuous distribution shift over time, as described in Gibbs and Candès, (2021).
@@ 3 Dynamic merging via exponential weighted majority
As illustrated in Section 1.1, our goal is to achieve a set that is as small as possible, particularly (7) asserts that this can be accomplished by learning a weight system that prioritizes methods yielding smaller sets. Furthermore, we want that the cumulative loss of our proposed method, indicated by $L_{M}^{(t)}:=\ell_{M}^{(1)}+\dots+\ell_{M}^{(t)}$, is not too large if compared to the cumulative loss incurred by the best expert up to time, indicated by
$L_{*}^{(t)}:=\min_{k\in[K]}L_{k}^{(t)}.$
@@
$w_{k}^{(1)}\leftarrow 1/K,L_{k}^{(0)}\leftarrow 0,$ $k=[K]$;
for rounds $t=1,\dots,T$ do
       $\mathcal{C}^{(t)}\leftarrow\left\{s\in\mathcal{S}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\{s\in\mathcal{C}_{k}^{(t)}\}>\frac{1+u^{(t)}}{2}\right\}$;
@@
Our proposed solution is to employ the AdaHedge algorithm, which dynamically adjusts the learning parameter over time as follows. The weights at each round are assigned according to $w^{(t)}_{k}\propto\exp\{-\eta^{(t)}L^{(t-1)}_{k}\}$, $k\in[K]$, with
@@
It is now possible to bound the loss of the sets produced by the majority vote method.
Theorem 3.1.
The loss of the weighted majority vote set over $T$ rounds can be bounded as
@@
As explained at the beginning, a relatively open question in the conformal prediction literature is how to do model selection or aggregation using conformal prediction. Since all methods do provide the same coverage level of $1-\alpha$ under i.i.d. data, we use a working definition of the “best” predictor as the one whose expected set dimension is smaller.
Below, we provide one answer to this question in an online data setting, by employing the algorithms introduced in the previous section. When our dynamic merging algorithm is used (with constant, adaptive or another stepsize rule) in this conformal context, we title the method as “conformal online model aggregation”, or COMA for short. In the i.i.d. setting, it is often the case that our adaptive stepsize update rule results in the weight vector rapidly putting almost full (unit) weight on the best predictor, resulting effectively in online model “selection”, but experiments in non-i.i.d. settings suggest that the weight vector can meaningfully fluctuate between different predictors if they perform better in different periods.
We start by examining our proposed methods, in a scenario where data are i.i.d., this implies that each algorithm at each iteration provides an interval with the guarantee described in (1). To examine the practical behavior of COMA, we apply it to a real-world dataset. Specifically, the dataset comprises 75 0007500075\,00075 000 observations pertaining to Airbnb apartments in New York City. The response variable is represented by the logarithm of the nightly price, while the covariates include information about the apartment and its geographical location. The data has been partitioned into 75 rounds, each consisting of 1 00010001\,0001 000 observations and a single test point. At each iteration, $K$ split-conformal prediction intervals based on the data observed at round $t$ are constructed using various algorithms. In the first simulation, $K=4$ regression algorithms were employed: linear model, lasso, ridge, and random forest. In the second scenario $K=5$ and also a neural network was utilized.
@@
The weights were adjusted using the COMA method with either fixed or adaptive $\eta$ settings as previously mentioned. For Algorithm 1 with fixed learning parameter throughout the iterations, two different values were selected, $\eta=0.1$ and $\eta=1$. In Figure 1, the hedge loss during various iterations in both scenarios is presented, along with the loss obtained by the best performing method(s). It can be seen that in both cases, that the hedge loss of the COMA algorithm with $\eta=1$ and of the adaptive COMA stabilizes at the loss level obtained by the random forest after a certain number of iterations, which is the best method. On the other hand, if $\eta$ is set to 0.1 the loss exhibits a slower decrease, attributed to the lower value of the learning parameter. In Figure 6 and Figure 5 of Appendix A, we report the weights assumed by the different regression algorithms during certain iterations. It can be seen that the COMA with adaptive parameter, after a few iterations, concentrates all the mass in the random forest, while the COMA algorithm with fixed learning parameter is more conservative in assigning the weights. This fact can be explained by the plot in Figure 7 of Appendix A where it is possible to note that in both cases ($K=4$ and $K=5$) the learning parameter $\eta^{(t)}$ is always higher than the levels $\eta=1$ and $\eta=0.1$.